### v1

# Multi-role Consensus through LLMs Discussions for Vulnerability Detection

Abstract
Recent advancements in large language models (LLMs) have highlighted the potential for vulnerability detection, a crucial component of software quality assurance. Despite this progress, most studies have been limited to the perspective of a single role, usually testers, lacking diverse viewpoints from different roles in a typical software development life-cycle, including both developers and testers. To this end, this paper introduces an approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code. Preliminary evaluation of the proposed approach indicates a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
Index Terms:
large language models; vulnerability detection; prompt engineering; software quality assurance
1 Introduction
As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats. Recently, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [1, 2].
However, a notable limitation in these studies is that they often only consider the dominant tester-centric perspective in the prompt, such as asking LLMs to act as quality assurance engineers to detect potential vulnerabilities. This single-role approach fails to capture the insights inherent in the real-world code review process, which typically benefits from the collaboration and discussion of team members (e.g., developer and tester) possessing diverse roles, responsibilities, and viewpoints.
To bridge this research gap, this paper introduces an approach, namely Multi-role Consensus through LLMs Discussions (MuCoLD), for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process. The key idea of MuCoLD is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
2 MuCoLD
Fig.1 shows the workflow of MuCoLD, where a tester and a developer in a code review team attempt to reach a collective consensus on the existence and classification of vulnerabilities in a code segment by discussing with each other according to their own perspectives through a three-stage process.
Initialization stage: The initiation stage is designed to enable the tester to independently provide an initial judgment. The tester receives the initial prompt detailing its role-setting, its task, and the code segment to analyze. And the tester is asked to output its first response, a textual completion that includes a judgment, restricted to a binary indicator (1 for vulnerable, 0 for non-vulnerable), and a brief reasoning. The response is constrained by a maximum token limit, ensuring that the tester’s reasoning is both precise and substantive, facilitating clarity and efficiency in the subsequent dialectic interaction. Then this initial judgment with reasoning is forwarded to the developer, together with a similar initial prompt.
Discussion stage: The discussion stage aims to realize an iterative output exchange in attempt to reach a collectively multi-perspective consensus inside the code review team. The tester and the developer, equipped with their unique perspective and initial judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities. During this stage, the tester and the developer repeats a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning. A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains goal-oriented and time-efficient.
Conclusion stage: The conclusion stage summarizes the discussions and output the final result. Once a consensus or the pre-set maximum discussion depth is reached, the tester’s latest judgment is recorded as the final judgment, as the tester usually holds primary responsibility in the review process.
Figure 1: Overview of MuCoLD
3 Preliminary Evaluation
The preliminary evaluation is driven by the research question as follows: In what ways does the proposed approach improve the performance of vulnerability detection?
3-A Experiment Settings
Dataset: The dataset used for the experiment, as referenced in [3], is a C/C++ dataset consisting of both vulnerable and non-vulnerable code segments across four categories: library/API function call (FC), arithmetic expression (AE), array usage (AU), and pointer usage (PU).
Task: LLMs are tasked with a binary classification on whether the provided code contains a major vulnerability that belongs to FC, AE, AU, or PU.
Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo model. MuCoLD is compared to the result obtained from the single-role LLM approach. Inspired by [1], the following two types of prompts are used to test these two approaches. The basic prompt directly asks the LLM about the presence of a major vulnerability in FC, AE, AU, or PU without additional context. Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [1, 4].
TABLE I: Experiment Results
Testing Vulner- Precision rate Recall rate F1 score
data ability single-role MuCoLD single-role MuCoLD single-role MuCoLD
[…]
Group1 FC 0.811 0.834 0.796 0.859 0.386 0.403 0.629 0.615 0.523 0.543 0.703 0.717
Group1 AE 0.791 0.787 0.794 0.828 0.298 0.384 0.623 0.626 0.432 0.516 0.698 0.713
Group1 AU 0.886 0.813 0.870 0.883 0.516 0.596 0.753 0.794 0.652 0.688 0.807 0.836
Group1 PU 0.829 0.778 0.829 0.830 0.430 0.511 0.689 0.743 0.566 0.617 0.752 0.784
Group2 FC 0.574 0.610 0.546 0.647 0.378 0.456 0.724 0.670 0.456 0.522 0.623 0.658
Group2 AE 0.561 0.503 0.534 0.622 0.332 0.364 0.684 0.656 0.417 0.422 0.599 0.639
Group2 AU 0.609 0.506 0.567 0.585 0.462 0.608 0.702 0.764 0.526 0.552 0.627 0.663
Group2 PU 0.601 0.523 0.547 0.551 0.466 0.540 0.698 0.824 0.525 0.531 0.613 0.660
Group3 FC 0.194 0.251 0.218 0.256 0.425 0.375 0.705 0.510 0.266 0.301 0.333 0.341
Group3 AE 0.239 0.234 0.236 0.258 0.330 0.410 0.665 0.535 0.277 0.298 0.348 0.349
Group3 AU 0.235 0.200 0.243 0.255 0.400 0.560 0.685 0.710 0.296 0.295 0.359 0.375
Group3 PU 0.219 0.205 0.222 0.232 0.450 0.525 0.695 0.790 0.295 0.295 0.336 0.359
1 Group1 consists 800 vulnerable and 200 non-vulnerable code segments; Group2: 500 and 500; Group3: 200 and 800.
3-B Results and Discussions
Table.I summarizes the experiment results, highlighting the improvements through MuCoLD. In average, there is a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score. In terms of computation costs, due to the need for conversation between different roles, it requires a 478% increase in the number of tokens consumed. The experiment results indicates a more obvious improvement in vulnerability detection performance, particularly when the proportion of vulnerable data in the testing dataset is higher. This is likely because the rounds of discussions allows LLMs to explore a broader spectrum of potential vulnerabilities, leading to a higher likelihood of classifying code as vulnerable. Moreover, the imprecise definitions tied to the vulnerability categories, each associated with a range of Common Weakness Enumeration IDs, may also lead more code segments identified as vulnerable.
4 Conclusion and Future Work
This paper proposes MuCoLD, an approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus. Preliminary evaluation have shown notable improvements in recall rates and F1 scores. Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, and extend its use to multi-class classification and vulnerability patching, broadening insights and proving its versatility.

### v2

@@
 Despite this progress, most studies have been limited to the perspective of a single role, usually testers, lacking diverse viewpoints from different roles in a typical software development life-cycle, including both developers and testers.
-To this end, this paper introduces an approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code.
+To this end, this paper introduces a multi-role approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code.
 Preliminary evaluation of the proposed approach indicates a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
@@
 1 Introduction
-As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats.
-Recently, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [1, 2].
+As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2].
+Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4, 5].
 However, a notable limitation in these studies is that they often only consider the dominant tester-centric perspective in the prompt, such as asking LLMs to act as quality assurance engineers to detect potential vulnerabilities.
 This single-role approach fails to capture the insights inherent in the real-world code review process, which typically benefits from the collaboration and discussion of team members (e.g., developer and tester) possessing diverse roles, responsibilities, and viewpoints.
-To bridge this research gap, this paper introduces an approach, namely Multi-role Consensus through LLMs Discussions (MuCoLD), for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process.
-The key idea of MuCoLD is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
-2 MuCoLD
-Fig.1 shows the workflow of MuCoLD, where a tester and a developer in a code review team attempt to reach a collective consensus on the existence and classification of vulnerabilities in a code segment by discussing with each other according to their own perspectives through a three-stage process.
+To bridge this research gap, this paper introduces a novel multi-role approach, namely Multi-role Consensus through LLMs Discussions, for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process.
+The key idea of this approach is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
+2 Multi-role Approach
 Initialization stage: The initiation stage is designed to enable the tester to independently provide an initial judgment.
@@
 Discussion stage: The discussion stage aims to realize an iterative output exchange in attempt to reach a collectively multi-perspective consensus inside the code review team.
-The tester and the developer, equipped with their unique perspective and initial judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities.
+The tester and the developer, equipped with their unique perspective and judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities.
 During this stage, the tester and the developer repeats a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning.
-A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains goal-oriented and time-efficient.
+A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains both goal-oriented and time-efficient.
 Conclusion stage: The conclusion stage summarizes the discussions and output the final result.
 Once a consensus or the pre-set maximum discussion depth is reached, the tester’s latest judgment is recorded as the final judgment, as the tester usually holds primary responsibility in the review process.
-Figure 1: Overview of MuCoLD
+Figure 1: Overview of the multi-role approach
 3 Preliminary Evaluation
@@
 Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo model.
-MuCoLD is compared to the result obtained from the single-role LLM approach.
+The proposed multi-role approach is compared to the result obtained from the single-role LLM approach.
 Inspired by [3], the following two types of prompts are used to test these two approaches.
@@
 Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
-Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [1, 4].
+Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [3].
 TABLE I: Experiment Results
 Testing Vulner- Precision rate Recall rate F1 score
-data ability single-role MuCoLD single-role MuCoLD single-role MuCoLD
+data ability single-role multi-role single-role multi-role single-role multi-role
 […]
@@
 3-B Results and Discussions
-Table.I summarizes the experiment results, highlighting the improvements through MuCoLD.
+Table.I summarizes the experiment results, highlighting the improvements through the proposed approach.
 In average, there is a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
@@
 4 Conclusion and Future Work
-This paper proposes MuCoLD, an approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus.
+This paper proposes a novel approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus.
 Preliminary evaluation have shown notable improvements in recall rates and F1 scores.
-Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, and extend its use to multi-class classification and vulnerability patching, broadening insights and proving its versatility.
+Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, broadening insights and proving its versatility.

### v3

@@
 Despite this progress, most studies have been limited to the perspective of a single role, usually testers, lacking diverse viewpoints from different roles in a typical software development life-cycle, including both developers and testers.
-To this end, this paper introduces a multi-role approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code.
-Preliminary evaluation of the proposed approach indicates a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
+To this end, this paper introduces a multi-role approach to employ LLMs to act as different roles simulating a real-life code review process and engaging in discussions toward a consensus on the existence and classification of vulnerabilities in the code.
+Preliminary evaluation of this approach indicates a 13.48% increase in the precision rate, an 18.25% increase in the recall rate, and a 16.13% increase in the F1 score.
 Index Terms:
@@
 As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2].
-Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4, 5].
+Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4].
 However, a notable limitation in these studies is that they often only consider the dominant tester-centric perspective in the prompt, such as asking LLMs to act as quality assurance engineers to detect potential vulnerabilities.
-This single-role approach fails to capture the insights inherent in the real-world code review process, which typically benefits from the collaboration and discussion of team members (e.g., developer and tester) possessing diverse roles, responsibilities, and viewpoints.
-To bridge this research gap, this paper introduces a novel multi-role approach, namely Multi-role Consensus through LLMs Discussions, for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process.
-The key idea of this approach is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
+This single-role approach results in an imperfect understanding of the intentions behind the code, an incomplete exploration of potential issues, and consequently, reduced detection efficiency.
+To avoid this, the real-world code review process usually involves the collaboration of team members (e.g., developer and tester) possessing diverse roles, responsibilities, and viewpoints.
+Inspired by this, this paper introduces a novel multi-role approach, namely Multi-role Consensus through LLMs Discussions, for employing LLMs to act as multiple different roles in vulnerability detection, simulating a real-life code review process.
+The key idea of this approach is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of the vulnerability in the code.
 2 Multi-role Approach
@@
 The tester receives the initial prompt detailing its role-setting, its task, and the code segment to analyze.
-And the tester is asked to output its first response, a textual completion that includes a judgment, restricted to a binary indicator (1 for vulnerable, 0 for non-vulnerable), and a brief reasoning.
+The tester is asked to output its first response, a textual completion that includes a judgment, restricted to a binary indicator (1 for vulnerable, 0 for non-vulnerable), and a brief reasoning.
 The response is constrained by a maximum token limit, ensuring that the tester’s reasoning is both precise and substantive, facilitating clarity and efficiency in the subsequent dialectic interaction.
 Then this initial judgment with reasoning is forwarded to the developer, together with a similar initial prompt.
-Discussion stage: The discussion stage aims to realize an iterative output exchange in attempt to reach a collectively multi-perspective consensus inside the code review team.
-The tester and the developer, equipped with their unique perspective and judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities.
-During this stage, the tester and the developer repeats a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning.
+Discussion stage: The discussion stage aims to realize an iterative output exchange in an attempt to reach a collectively multi-perspective consensus inside the code review team.
+The tester and the developer, equipped with their unique perspective and judgments, enter a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities.
+During this stage, the tester and the developer repeat a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning.
 A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains both goal-oriented and time-efficient.
-Conclusion stage: The conclusion stage summarizes the discussions and output the final result.
+Conclusion stage: The conclusion stage summarizes the discussions and outputs the final result.
 Once a consensus or the pre-set maximum discussion depth is reached, the tester’s latest judgment is recorded as the final judgment, as the tester usually holds primary responsibility in the review process.
@@
 Dataset: The dataset used for the experiment, as referenced in [5], is a C/C++ dataset consisting of both vulnerable and non-vulnerable code segments across four categories: library/API function call (FC), arithmetic expression (AE), array usage (AU), and pointer usage (PU).
-Task: LLMs are tasked with a binary classification on whether the provided code contains a major vulnerability that belongs to FC, AE, AU, or PU.
-Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo model.
+Task: LLMs are tasked with a binary classification on whether the provided code segment contains any vulnerability that belongs to FC, AE, AU, or PU.
+Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo-0125 model.
 The proposed multi-role approach is compared to the result obtained from the single-role LLM approach.
 Inspired by [3], the following two types of prompts are used to test these two approaches.
-The basic prompt directly asks the LLM about the presence of a major vulnerability in FC, AE, AU, or PU without additional context.
+The basic prompt directly asks the LLM about the presence of any vulnerability in FC, AE, AU, or PU without additional context.
 Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
@@
 data ability single-role multi-role single-role multi-role single-role multi-role
-[…]
-Group1 FC 0.811 0.834 0.796 0.859 0.386 0.403 0.629 0.615 0.523 0.543 0.703 0.717
-Group1 AE 0.791 0.787 0.794 0.828 0.298 0.384 0.623 0.626 0.432 0.516 0.698 0.713
-Group1 AU 0.886 0.813 0.870 0.883 0.516 0.596 0.753 0.794 0.652 0.688 0.807 0.836
-Group1 PU 0.829 0.778 0.829 0.830 0.430 0.511 0.689 0.743 0.566 0.617 0.752 0.784
-Group2 FC 0.574 0.610 0.546 0.647 0.378 0.456 0.724 0.670 0.456 0.522 0.623 0.658
-Group2 AE 0.561 0.503 0.534 0.622 0.332 0.364 0.684 0.656 0.417 0.422 0.599 0.639
-Group2 AU 0.609 0.506 0.567 0.585 0.462 0.608 0.702 0.764 0.526 0.552 0.627 0.663
-Group2 PU 0.601 0.523 0.547 0.551 0.466 0.540 0.698 0.824 0.525 0.531 0.613 0.660
-Group3 FC 0.194 0.251 0.218 0.256 0.425 0.375 0.705 0.510 0.266 0.301 0.333 0.341
-Group3 AE 0.239 0.234 0.236 0.258 0.330 0.410 0.665 0.535 0.277 0.298 0.348 0.349
-Group3 AU 0.235 0.200 0.243 0.255 0.400 0.560 0.685 0.710 0.296 0.295 0.359 0.375
-Group3 PU 0.219 0.205 0.222 0.232 0.450 0.525 0.695 0.790 0.295 0.295 0.336 0.359
+group1 category basic CoT basic CoT basic CoT basic CoT basic CoT basic CoT
+Group1 FC 0.735 0.756 0.830 0.837 0.610 0.643 0.801 0.796 0.667 0.695 0.816 0.816
+Group1 AE 0.750 0.756 0.837 0.854 0.618 0.634 0.779 0.818 0.677 0.689 0.807 0.835
+Group1 AU 0.772 0.759 0.860 0.856 0.668 0.695 0.851 0.896 0.716 0.725 0.856 0.875
+Group1 PU 0.753 0.769 0.832 0.847 0.629 0.654 0.791 0.875 0.685 0.707 0.811 0.861
+Group2 FC 0.568 0.564 0.637 0.641 0.640 0.660 0.752 0.718 0.602 0.608 0.690 0.677
+Group2 AE 0.575 0.553 0.627 0.647 0.628 0.640 0.732 0.750 0.600 0.593 0.675 0.694
+Group2 AU 0.549 0.579 0.651 0.639 0.652 0.696 0.750 0.778 0.596 0.632 0.697 0.702
+Group2 PU 0.550 0.536 0.619 0.632 0.674 0.698 0.778 0.782 0.606 0.606 0.690 0.699
+Group3 FC 0.196 0.204 0.225 0.232 0.650 0.695 0.730 0.765 0.302 0.315 0.344 0.356
+Group3 AE 0.197 0.198 0.229 0.229 0.660 0.680 0.755 0.765 0.304 0.306 0.352 0.352
+Group3 AU 0.218 0.211 0.238 0.246 0.735 0.710 0.805 0.830 0.336 0.325 0.368 0.380
+Group3 PU 0.199 0.197 0.224 0.237 0.690 0.705 0.755 0.790 0.308 0.309 0.345 0.364
 1 Group1 consists 800 vulnerable and 200 non-vulnerable code segments; Group2: 500 and 500; Group3: 200 and 800.
 3-B Results and Discussions
-Table.I summarizes the experiment results, highlighting the improvements through the proposed approach.
-In average, there is a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
-In terms of computation costs, due to the need for conversation between different roles, it requires a 478% increase in the number of tokens consumed.
-The experiment results indicates a more obvious improvement in vulnerability detection performance, particularly when the proportion of vulnerable data in the testing dataset is higher.
-This is likely because the rounds of discussions allows LLMs to explore a broader spectrum of potential vulnerabilities, leading to a higher likelihood of classifying code as vulnerable.
-Moreover, the imprecise definitions tied to the vulnerability categories, each associated with a range of Common Weakness Enumeration IDs, may also lead more code segments identified as vulnerable.
+Table I summarizes the experiment results, highlighting the improvements through the proposed approach.
+On average, there is a 13.48% increase in the precision rate, an 18.25% increase in the recall rate, and a 16.13% increase in the F1 score.
+In terms of computation costs, due to the need for conversation between different roles, it requires a 484% increase in the number of tokens consumed.
+An illustrative discussion record demonstrating how the developer contributed to enhancing the tester’s detection accuracy is as follows.
+Initially, the tester identified a code segment as vulnerable due to the lack of validation or sanitization during the concatenation of two buffers.
+However, as the developer pointed out that the wcsncat function and the proper size argument are used to prevent a buffer overflow, the tester finally made the correct judgment that this code segment is non-vulnerable.
+Given the limitations on paper length, please refer to the complete discussion record at github.com/rockmao45/LLMVulnDetection.
+The experiment results indicate a more obvious improvement, especially in the recall rate and F1 score when the proportion of vulnerable data in the testing dataset is higher.
+This is likely because the rounds of discussions allows LLMs to explore a broader spectrum of potential vulnerabilities.
 4 Conclusion and Future Work
-This paper proposes a novel approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus.
-Preliminary evaluation have shown notable improvements in recall rates and F1 scores.
-Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, broadening insights and proving its versatility.
+This paper proposes an approach for improving vulnerability detection with LLMs, where LLMs act as different roles in a real-life code review team, discussing the existence and classification of vulnerabilities towards a consensus.
+Preliminary evaluations have shown notable improvements in recall rates and F1 scores.
+Future work should integrate in-context learning to guide LLMs to better collaborate in their discussions.

### v4

@@
 As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2].
-Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4].
+Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4, 5].
 However, a notable limitation in these studies is that they often only consider the dominant tester-centric perspective in the prompt, such as asking LLMs to act as quality assurance engineers to detect potential vulnerabilities.
@@
 Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
+Discussion constraints: The maximum discussion depth is set to 5, and the maximum response length is set to 120 tokens.
 Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [3].
@@
 4 Conclusion and Future Work
-This paper proposes an approach for improving vulnerability detection with LLMs, where LLMs act as different roles in a real-life code review team, discussing the existence and classification of vulnerabilities towards a consensus.
-Preliminary evaluations have shown notable improvements in recall rates and F1 scores.
-Future work should integrate in-context learning to guide LLMs to better collaborate in their discussions.
+This paper proposes a novel approach for improving vulnerability detection with LLMs, where LLMs act as different roles in a real-life code review team, discussing the existence and classification of vulnerabilities towards a consensus.
+Preliminary evaluations have shown notable improvements in precision rates, recall rates and F1 scores.
+Future work should enhance this approach by integrating in-context learning to guide LLMs to better collaborate in their discussions.